# Uncovering Vulnerabilities of LLM-Assisted Cyber Threat Intelligence

### v2

@@ 3.2 Autoregressive failure mode determination (RQ2)
3.3 Human-in-the-loop categorization of failure instances (RQ3)
@@
Step 1 (Multi-agent decision). For each threat instance $x_{i}$ requiring failure mode classification, we construct a model set $\Theta=\{\texttt{GPT-5},\texttt{Llama-4-17B},\texttt{Gemini-2.5},\texttt{Claude-Sonnet-4}\}$. Each model independently proposes an initial label $\hat{y}_{i}^{(1)}\in\mathcal{T}$ during the first execution.
Step 2 (Repetition for stability). Motivated by evidence that unstable internal knowledge may lead to fluctuations across runs (Kumar et al., 2024), we execute a second round of multi-agent deliberation. In this round, each LLM observes the predictions $\hat{y}_{i}^{(1)}$ from other models, refines its reasoning, and then proposes a new label $\hat{y}_{i}^{(2)}$ for the same instance.
@@
The multi-agent collaboration handles the majority of straightforward cases efficiently, while humans focus only on instances with instability or misalignment (empirically, less than 1.8% cases in this step), thus balancing scalability with reliability in failure mode categorization (Algorithm in 3).
4 Research Finding
Our categorization method leads to a detailed list of vulnerability types, as presented in Figure 2. Building on this, we further investigate specific failure cases to address the following questions:
RQ4: How do different LLMs present varying proportions of vulnerabilities?
RQ5: What are the primary root causes that give rise to these vulnerabilities?
RQ6: How do multiple vulnerabilities interact and become intertwined within the same instance?
4.1 Varying Performance of LLMs (RQ4)
@@
We first study the varying performance of different models. As shown in Figure 4, and consistent with the distributions across vulnerabilities in Figure 2, our analysis reveals clear differences in how various LLMs handle distinct categories of errors: For vulnerabilities that directly impair LLM inference (e.g., co-mentioned but irrelevant mitigation strategies or outdated metadata that reduce retrieval effectiveness), general-purpose LLMs tend to accumulate a higher volume of failures. This indicates that their broad but non-specialized training leaves them vulnerable to overfitting and misinterpreting contextual evidence.
In contrast, for vulnerabilities rooted in the data used during fine-tuning (e.g., confounding factors in the threat corpus), cyber agents such as Foundation-Sec or Cyber-Zero tend to produce less reliable outputs (e.g., forecasting exploitability with contradictory PoCs). Such vulnerabilities act as data poisoning, where spurious correlations and contradictory knowledge can arise either intentionally (introduced by adversaries) or unintentionally (inherent in the fragmented cyber threat landscape).
4.2 Root Causes of Vulnerabilities (RQ5)
@@
During failure categorization (§3), we also gained detailed insights into how different vulnerabilities are triggered by the used techniques among CTI tasks. Table 3 summarizes these causes in alignment with the vulnerability types shown in Figure 2. Here, we highlight some representative cases:
Co-mention bias (1.1) in ❶ contextualization. Co-mention bias emerges in contextualization when retrieval systems treat co-occurring entities in raw reports as causally linked. The security bulletin may list multiple vulnerabilities or mitigation strategies together in a single section, even though only one is relevant to a particular incident. Topic modeling or knowledge base mapping that lack fine-grained data disambiguation may surface all co-mentioned entities as equally relevant. This inflates the context with spurious links (e.g., assigning unrelated CVEs or MITRE TTPs to the same intrusion set), misguiding LLMs to propagate spurious associations.
@@
Case Study. Microsoft Patch Tuesday advisories often list many CVEs under one product (e.g., Windows Server) (Microsoft Security Response Center, 2003). The contextualization pipeline ingests the bulletin and a threat report cites CVE-2021-34527 (PrintNightmare), GPT-5 incorrectly infers that all co-mentioned CVEs were exploited in the same campaign, creating false links between benign vulnerabilities and active threats.
Conflicting report (2.2) in ❹ Mitigation. Mitigation is influenced by different testing environment and study scope of different vendors. For example, one advisory may claim that a patch fully resolves an exploited vulnerability, while another report provides evidence that attackers continued exploitation through chained dependencies, such as leveraging an adjacent misconfiguration. Mitigation mapping that relies on these contradictory reports may overestimate or underestimate the coverage of a particular countermeasure. Similarly, summarization pipelines may output inconsistent defensive playbooks, some emphasizing patch deployment while others prioritize compensating controls. These varying mitigation strategies leaving LLMs uncertain about actions that truly address threat incidents.
@@
Case Study. In the 2021 Microsoft Exchange “Hafnium” case (Microsoft Threat Intelligence & Microsoft 365 Security, 2021), some advisories claimed patches fully mitigated the threat, while others warned of persistent web shells post-patch. The conflicting reports gave Qwen retriever false assurance about their mitigation effectiveness.
@@
A comprehensive analysis with case studies across all vulnerabilities is provided in Appendix D.1.
4.3 Study of Intertwined Vulnerabilities (RQ6)
@@
We also notice that multiple vulnerabilities frequently interact and compound within the same instance. The root cause lies in the coupled nature of the CTI pipeline: early retrieval errors such as co-mention bias or skewed sources propagate forward as “facts,” which downstream attribution or prediction models cannot easily disconfirm. This effect is amplified by heterogeneous and drifting evidence, wherein reports that differ temporally, semantically, or structurally force models to merge incompatible signals. Under such uncertainty, models lean on inductive shortcuts, such as memorized actor–TTP links, which transform incomplete or conflicting metadata into brittle reasoning. Finally, the hierarchical and dependency-driven structure of attack chains and mitigation mappings means that once one link is misread, the error cascades into adjacent stages. As a result, intertwined vulnerabilities like co-mention bias plus temporal contradiction, or unseen patterns combined with environmental unawareness, emerge as self-reinforcing loops that entangle multiple failures within the same CTI instance.

### v3

@@ 3.2 Failure model determination (RQ2)
3.3 Human-in-the-loop decision (RQ3)
@@
Step 1 (Multi-agent decision). For each threat instance $x_{i}$ requiring failure mode classification, we construct a model set $\Theta=\{\text{GPT-5},\text{Llama-4-17B},\text{Gemini-2.5},\text{Claude-Sonnet-4}\}$. Each model independently proposes an initial label $\hat{y}_{i}^{(1)}\in\mathcal{T}$ during the first execution.
Step 2 (Repetition for stability). Motivated by evidence that unstable internal knowledge may lead to fluctuations across runs (kumar2024training), we execute a second round of multi-agent deliberation. In this round, each LLM observes the predictions $\hat{y}_{i}^{(1)}$ from other models, refines its reasoning, and then proposes a new label $\hat{y}_{i}^{(2)}$ for the same instance.
@@
The multi-agent collaboration efficiently handles most of routine cases , while humans focus only on instances with instability or misalignment (empirically, less than 1.8% cases in this step), thus balancing scalability with reliability in failure mode categorization (Algorithm in 3).
4 Experimental Result
Our categorization method leads to a detailed analysis of vulnerability types, as presented in Figure 2. Building on this, we further investigate specific failure cases and potential defense strategies to address the following questions:
Q1: Is the proposed method effective in categorizing vulnerabilities within the threat landscape?
Q2: How do different components of the proposed method influence its overall effectiveness?
Q3: How do different LLMs exhibit varying categorizations?
Q4: What are the root causes of these vulnerabilities?
Next, we conduct detailed experiments to answer these questions using the same settings and models as in §2.2. We further detail our defensive solutions in §5.
4.1 Effectiveness of vulnerability categorization with baselines (Q1)
Baseline setting. To validate the efficacy of our proposed framework, we benchmark it against several basic methodologies: (i) Clustering aggregates evaluation results of CTI benchmarks (§2.2) into unsupervised groups based on vector embedding similarity using K-Means. (ii) Topic modeling leverages latent dirichlet allocation to statistically extract semantic structures and keywords from the benchmarking results (§2.2). (iii) GPT judge, wherein GPT-5 independently classifies failure modes based on provided definitions. (iv) Gemini judge similarly leverages the Gemini model to categorize vulnerabilities. (v) Multi-agent debate constructs an iterative consensus mechanism where multiple LLM agents (same as in §3.3) refine each other’s classifications to reduce individual hallucinations.
Results. The comparison shows clear distributional gaps between unsupervised baselines and semantic reasoning approaches. Statistical methods such as clustering and topic modeling miss many of the fine-grained subtype boundaries, while producing a skewed failure distribution that departs substantially from the human reference. LLM-based judges align more closely with human judgments, but still exhibit drift on the more challenging categories that require multi-step reasoning. Overall, our method yields a subtype distribution that is consistently closest to human annotation, correcting the systematic skews seen in both single-agent predictions and simple consensus-style debates.
Explanations. The remaining gap largely reflects how well each method handles context-dependent ambiguity. Topic modeling, for example, struggles with Confounding factors (1.3) and Temporal contradictions (2.1) because these errors are driven by latent causal structure and logical inconsistency rather than surface-level lexical similarity. Standard LLM judges reduce this mismatch, but they often over-attribute cases to Spurious correlations (1.x) and inflate Overfitted reasoning (3.3), effectively “finding” patterns that are plausible in the abstract yet unsupported by the specific incident context. Our approach mitigates these tendencies by grounding decisions in explicitly retrieved evidence and structured multi-agent deliberation, which improves attribution for high-nuance categories such as Conflicting reports (2.2) and Unseen patterns (3.2) that other methods frequently misclassify.
Table 4: Ablation study by removing components from proposed vulnerability categorization method.
Method Spurious Correlation Contradictory Knowledge Constrained Generalization
1.1 1.2 1.3 1.4 1.5 2.1 2.2 2.3 2.4 2.5 2.6 3.1 3.2 3.3 3.4
Co-men. Exploit. Confond. Skewed Hier. Temporal Conflict Semantic Diver. Misal. Counter. Distribut. Unseen Overfit. Env.
bias bias factor source metadata contra. report conflict data knowl. CTI bias pattern reasoning unaware
Human annotation 4.14 1.71 15.72 6.16 2.05 13.98 22.29 4.14 4.53 1.97 7.45 21.79 23.13 7.64 11.76
w/o stratification 12.62 5.14 10.37 11.23 5.89 17.45 16.12 9.87 8.45 5.23 12.56 17.89 16.34 14.56 15.23
w/o human initiation 4.93 2.56 14.23 7.82 3.12 12.45 20.56 5.45 5.78 2.89 9.12 20.34 21.67 11.45 12.34
w/o multi-agent decision 2.34 6.78 21.56 3.45 0.56 9.12 28.45 1.67 2.89 0.92 4.12 30.23 13.45 5.67 6.78
w/o repetition 9.45 6.23 11.67 10.12 6.34 19.56 17.89 9.12 7.56 6.12 13.45 19.67 17.23 15.45 14.56
w/o human verification 3.67 2.89 16.45 5.45 2.67 14.89 23.67 4.89 3.78 2.56 6.45 23.12 24.56 9.34 10.23
Ours 4.12 1.87 15.50 6.31 1.98 13.86 22.33 3.96 4.49 1.85 7.77 23.34 25.06 10.02 10.80
@@ 4.1 Effectiveness of vulnerability categorization with baselines (Q1)
4.2 Ablation study (Q2)
Results. The ablation study shows a clear dependency across components. When we remove core architectural modules, especially multi-agent debate and stratification, the predicted subtype distribution departs sharply from the human gold standard, with unstable spikes emerging in categories such as Confounding Factors (1.3) and Overfitted Reasoning (3.3). By comparison, disabling the human-in-the-loop steps (human initiation and verification) causes a smaller but still noticeable drop: the distribution drifts away from the ground truth, yet it remains far more stable than in settings where the automated reasoning machinery is removed.
Explanations and efficiency analysis. These results highlight a primary strength of our framework: it cuts manual workload substantially without sacrificing efficacy. Notably, even the w/o human verification variant still outperforms fully automated baselines (e.g., Clustering and Topic Modeling), indicating that the multi-agent architecture is doing most of the cognitive heavy lifting for failure categorization. In practice, this reframes the analyst’s role from labor-intensive labeling to lightweight oversight—spot-checking and validating outputs rather than building labels from scratch. As a result, CTI teams can approach near-human precision while spending only a fraction of the time typically required for manual classification.
@@ 4.2 Ablation study (Q2)
Case Study I: co-mention bias (Type 1.1) in ❶ contextualization Microsoft Patch Tuesday advisories often list many CVEs under one product (e.g., Windows Server) (microsoft_patch_tuesday). The contextualization pipeline ingests the bulletin and a threat report cites CVE-2021-34527 (PrintNightmare), GPT-5 incorrectly infers that all co-mentioned CVEs were exploited in the same campaign, creating false links between irrelevant CVEs and active threats.
@@
Case Study II: conflicting report (Type 2.2) in ❹ mitigation In the 2021 Microsoft Exchange “Hafnium” case (MicrosoftHafnium2021), some advisories claimed patches fully mitigated the threat, while others warned of persistent web shells post-patch. The conflicting reports gave Qwen retriever false assurance about their mitigation effectiveness.
4.3 Varying LLM performance (Q3)
Results and explanations. We further study the varying performance of different models. As shown in Figure 4, and consistent with the distributions across vulnerabilities in Figure 2, our analysis reveals clear differences in how various LLMs handle distinct categories of errors: For vulnerabilities that directly impair LLM inference (e.g., co-mentioned but irrelevant mitigation strategies or outdated metadata that reduce retrieval effectiveness), general-purpose LLMs tend to accumulate a higher volume of failures. This indicates that their broad but non-specialized training leaves them vulnerable to overfitting and misinterpreting contextual evidence.
In contrast, for vulnerabilities rooted in the fine-tuning data (e.g., confounding factors in the threat corpus), cyber agents such as Foundation-Sec or Cyber-Zero tend to produce less reliable outputs (e.g., forecasting exploitability with contradictory PoCs). Such vulnerabilities act as data poisoning, where spurious correlations and contradictory knowledge can arise either intentionally (introduced by adversaries) or unintentionally (inherent in the fragmented cyber threat landscape).
Case Study III: The intertwined threats – MOVEit (CVE-2023-34362) Early CTI reporting showed contextualization failure (Type 1.3): triage lumped the SQL injection into generic IIS exploitation, and co-mention bias hid the specific bug in moveitisapi.dll. Later, Attribution failure (Type 2.1) misattributed the activity to Silence due to the "Truebot" downloader, despite later evidence pointing solely to Cl0p (Lace Tempest). Finally, Prediction failure (Type 3.2) underestimated the Impact Forecast, missing the supply-chain cascade (e.g., Zellis, BBC) because the model overgeneralized from typical single-victim ransomware cases.
4.4 Root cause analysis (Q4)
Table 6: Absolute improvement ($\Delta$) in CTI task performance following causal interventions. The proposed method targets all three failure types (spurious correlation, contradictory knowledge, and constrained Generalization) at inference time reasoning. Each number denotes the raw gain (e.g., F1 + $\Delta$) over the original performance (Table 2).
Detailed CTI Task Metric Industry-leading, general-purpose Cybersecurity-specialized
G5 Go4 CLD GEM LL70 MIX QWN GRK FSC CB0 ZYS LLY CBS SPT DHT
❶ Contextualization
Affected Systems F1 .031 .042 .055 .068 .022 .061 .054 .041 .038 .041 .035 .033 .039 .042 .038
Attack Infrastructure F1 .025 .048 .062 .071 .031 .068 .065 .059 .052 .055 .041 .038 .045 .048 .044
Vulnerability Linking Acc .042 .051 .068 .075 .048 .072 .068 .062 .058 .061 .039 .035 .041 .044 .039
Malware Family Mapping F1 .045 .058 .071 .082 .055 .079 .074 .068 .062 .065 .048 .042 .051 .055 .049
IOC Normalization F1 .018 .025 .035 .041 .022 .038 .035 .031 .029 .032 .021 .018 .025 .028 .022
Threat Report Alignment BLEU .082 .095 .112 .098 .075 .105 .098 .089 .091 .085 .062 .058 .075 .088 .065
Event Timeline Construction BLEU .075 .088 .095 .091 .082 .094 .089 .081 .085 .082 .055 .048 .058 .062 .055
Graph Population Acc .035 .052 .075 .081 .042 .085 .072 .088 .068 .075 .045 .042 .038 .065 .035
Source Reliability Scoring AUC .012 .015 .028 .021 .018 .032 .025 .022 .025 .028 .015 .012 .018 .022 .015
❷ Attribution
Threat Actor Linking Acc .021 .028 .055 .042 .035 .052 .048 .065 .058 .062 .025 .018 .028 .025 .035
TTP Extraction F1 .032 .038 .048 .052 .045 .055 .051 .048 .042 .045 .031 .028 .035 .038 .032
Campaign Attribution Acc .048 .055 .068 .072 .052 .075 .071 .065 .058 .061 .042 .038 .045 .048 .042
Infrastructure Reuse F1 .052 .061 .075 .081 .058 .082 .078 .072 .065 .068 .045 .041 .048 .052 .045
Language/Style Profiling Acc .025 .032 .045 .048 .038 .052 .048 .045 .041 .044 .028 .025 .032 .035 .028
False Flag Detection F1 .088 .095 .108 .112 .082 .115 .105 .098 .095 .092 .075 .068 .082 .085 .078
Evidence Weighting BLEU .092 .105 .125 .115 .088 .118 .112 .105 .098 .095 .072 .065 .085 .092 .075
Relation Graph Building F1 .042 .051 .062 .068 .048 .065 .061 .058 .052 .055 .038 .035 .042 .045 .038
❸ Prediction
Exploit Likelihood AUC .015 .022 .035 .038 .018 .042 .038 .032 .045 .028 .022 .018 .015 .025 .018
Impact Forecast BLEU .065 .078 .092 .085 .058 .095 .088 .082 .075 .072 .052 .048 .058 .065 .052
Target Sector Prediction Acc .028 .035 .052 .045 .032 .055 .051 .058 .048 .042 .035 .031 .028 .038 .032
Campaign Escalation AUC .045 .058 .072 .078 .052 .075 .071 .068 .062 .065 .048 .045 .052 .055 .048
❹ Mitigation
Patch Recommendation F1 .038 .045 .058 .062 .032 .065 .061 .055 .052 .055 .041 .038 .032 .035 .045
Rule Generation (YARA) BLEU .052 .065 .082 .075 .058 .085 .078 .072 .068 .065 .045 .042 .052 .055 .048
Response Summarization BLEU .042 .055 .068 .062 .048 .072 .065 .058 .055 .052 .038 .035 .032 .045 .038
Mitigation–TTP Mapping Acc .035 .042 .055 .058 .041 .062 .058 .052 .048 .051 .032 .028 .031 .035 .032
Defensive Playbook Gen BLEU .062 .075 .088 .082 .068 .092 .085 .078 .072 .068 .055 .051 .052 .058 .055
Countermeasure Ranking NDCG .055 .068 .082 .078 .061 .085 .078 .072 .065 .068 .048 .045 .042 .055 .058
Incident Ticket Generation Acc .025 .032 .045 .038 .018 .048 .045 .042 .038 .045 .028 .025 .022 .028 .025
We conduct vulnerability-specific inspections across different CTI stages, investigating why a vulnerability may arise in a given stage (or why it may not) (§4.4.1 and App.  D.1). Besides, we observe that vulnerabilities can be intertwined within the same CTI stage; we provide a detailed analysis in §4.4.2 and App.  D.2.
4.4.1 Analyzing vulnerabilities within a single CTI stage
During failure categorization (§3), we gained insight into how different vulnerabilities are triggered by the techniques used across CTI tasks. Table 5 summarizes these drivers, aligned with the vulnerability types in Figure 2. Below, we highlight representative cases.
Co-mention bias (Type 1.1) in ❶ contextualization. As shown in Case Study I, co-mention bias arises during contextualization when retrieval systems treat entities that co-occur in raw reports as if they were causally related. Security bulletins often list multiple vulnerabilities, indicators, or mitigations in a single section, even though only a subset is relevant to the incident at hand. Methods such as topic modeling or coarse knowledge-base mapping, which lack fine-grained disambiguation, may surface all co-mentioned entities as equally relevant. This bloats the retrieved context with spurious links (e.g., attaching unrelated CVEs or MITRE TTPs to the same intrusion), which in turn prompts LLMs to propagate incorrect associations.
Conflicting reports (Type 2.2) in ❹ mitigation. As shown in Case Study II, mitigation guidance often diverges across vendors due to differences in testing environments, threat visibility, and evaluation scope. For example, one advisory may claim a patch fully addresses an exploited vulnerability, while another provides evidence of continued exploitation via chained dependencies (e.g., an adjacent misconfiguration). Mitigation mapping built on such contradictory sources can overestimate or underestimate the effectiveness of a countermeasure. Similarly, summarization pipelines may produce inconsistent playbooks (e.g., some prioritizing patch deployment and others emphasizing compensating controls) leaving LLMs uncertain about actions truly mitigate the incident.
Semantic conflict (Type 2.3) in ❷ attribution (less influential). Semantic conflict is less influential in attribution because actor identification depends on relational evidence and behavioral signatures. In practice, attribution relies more on infrastructure reuse, campaign TTPs, and stylistic cues than on whether a malware family is labeled “PlugX” or “Korplug.” Relation extraction and event-graph construction emphasize these structural patterns, making attribution models relatively tolerant to naming inconsistencies as long as the underlying relationships remain stable. While semantic divergence can introduce minor noise, it rarely flips the actor assignment: different labels for the same family typically still connect to the same infrastructure nodes and TTPs. As a result, attribution is comparatively robust to semantic conflicts.
@@ 4.4.1 Analyzing vulnerabilities within a single CTI stage
A comprehensive analysis with case studies across all vulnerability types is provided in Appendix D.1.
4.4.2 Analyzing intertwined vulnerabilities
@@
As shown in Case Study III, we also observe that multiple vulnerabilities may interact within the same CTI instance. This stems from the tightly coupled nature of CTI pipelines: early-stage retrieval issues (e.g., co-mention bias or source skew) are carried forward as “facts,” leaving downstream attribution and prediction models with little room to falsify them. The problem is further amplified by heterogeneous and drifting evidence that force LLMs to memorize signals that are not truly compatible. Under this uncertainty, models fall back on inductive shortcuts (e.g., memorized actor-TTP associations), turning partial or noisy metadata into brittle reasoning. Finally, because CTI reasoning is hierarchical and dependency-driven (e.g., attack chains and mitigation mappings), misreading one link can cascade into adjacent stages. As a result, vulnerability combinations such as co-mention bias paired with temporal contradiction, or unseen patterns coupled with environmental unawareness, can form self-reinforcing loops that entangle multiple failures within a single instance.